40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Sun Jun 29 02:33:28 2008, Koji, Update, General, Abs. Len. Meas. ~ Optical setup (I) optical_configuration.pngoptical_layout_ap_table2.pngoptical_layout_ap_table2.pdf
    Reply  Mon Jun 30 05:33:38 2008, Koji, Update, General, Abs. Len. Meas. ~ Optical setup (II) 
       Reply  Tue Jul 1 21:54:05 2008, Koji, Update, General, Re: Abs. Length Meas. setup 
    Reply  Wed Jul 2 15:14:42 2008, steve, Update, General, added beam traps 
Message ID: 599     Entry time: Mon Jun 30 05:33:38 2008     In reply to: 590     Reply to this: 619
Author: Koji 
Type: Update 
Category: General 
Subject: Abs. Len. Meas. ~ Optical setup (II) 
o The position of the iris was adjusted so as not to disturub the beam for OMCR CCD.

o The RF spectrum analyzer was returned to the place of the network analyzer.


Quote:

In the process of making this report, I noticed that one of the iris apertures is about disturbing the beam for OMCR CCD. I will check this before I go to Hanford. Also an RF spectrum analyzer is at the AP table. I try to return this near the PSL on Monday morning.
ELOG V3.1.3-